A washer and a dryer cost $601 combined. The washer costs $49 less than the dryer. What is the cost of the dryer?

I came up with $276, but I don't know if it is right..please help thanks

Found 3 solutions by Fombitz, checkley77, gonzo:
Answer by Fombitz(32388)

No, you've got it backwards.
Check your variables again.
The washer should cost $276 since it costs less than the dryer.

Answer by checkley77(12844)

X+(X-49)=601
2X=601+49
2X=650
X=650/2
X=325 FOR THE DRYER PRICE.
601-325=276 FOR THE WASHER PRICE.


Answer by gonzo(654)

w = cost of washer, d = cost of dryer.
w + d = 601 dollars.
w = d - 49 (washer cost 49 dollars less than the cost of the dryer).
substituting in original equation and solving for the cost of the dryer .............
d - 49 + d = 601
2d = 601 + 49.
2d = 650.
d = 650 / 2 = 325 dollars.
so the dryer cost 325 dollars and the washer cost 49 dollars less than the dryer = 276 dollars.
325 + 276 = 601.
it looks like you came up with the cost of the washer.